Inverse operations is a concept that needs to be understood because it is often used in algebra to solve variables.

We explain inverse operations as applied in numerical statements. Inspect the following example of a numerical equation, which is later translated into an algebraic equation:

21 × 5 - 10 = 95

21 × 5 = 95 + 10 (Add 10 both sides of =.)

21 = (Divide by 5 both sides of =.)

   Do the following:

Following our example to the left, work through the numerical equation below and then translate it into an algebraic equation, where you solve for the unknown value:

3/5 + 13 = 20

 


1
Expert's answer
2021-05-24T17:20:20-0400

"\frac{3}{5}x+13=20"

"\frac{3}{5}x+13-13=20-13"

"\frac{3}{5}x=7"

"x=7\times\frac{5}{3}"

"x=\frac{35}{3}"

"x=11.67"